OpenStudy (anonymous):

4| 2s + 5 | = 8 solve and check your solution

OpenStudy (john_es):

If (2s+5)<0, that happens when s<-5/2, then \[4(-2s-5)=8\Rightarrow -8s-20=8\Rightarrow s=-28/8=-7/2\] If (2s+5)>0, that happens, when s>-5/2, then, \[4(2s+5)=8\Rightarrow 8s+20=8\Rightarrow s=-12/8=-3/2\] So you have 2 solutions. To chek the solutions, simply put the obtained values in the original equation of the problem.